Name: Andrew Marzo

Content and Structure – The paper

1. Has a clearly stated topic or question that it will address – 10

2. Does address the topic or question – 5

3. Has a clearly articulated and logical structure:

a. First paragraph lays out the topic or question – 10 b. Later paragraphs have topic sentences and clearly identifiable main ideas – 10 c. Transitions between paragraphs are logical – 10

4. Accurately presents facts and viewpoints of sources – 10

5. Provides sufficient textual evidence to support the argument – 10

6. Is concise, clear, and not too repetitive – 8

7. Has a clearly identifiable conclusion – 10

8. Has a convincing argument that draws together the evidence presented – 10

Total: 93 (A-)

Style and English Usage – The paper

1. Is clear, direct, and appropriate in tone – 20

2. Is free of misspellings – 10

3. Is free of grammatical mistakes – 14

4. Avoids incomplete sentences – 10

5. Avoids run-on or overly complicated sentences – 9

6. Uses correct punctuation – 8

7. Is free of jargon, cliches, and misused words – 10

8. Cites quotations and/or references correctly – 10

Total: 96 (A)

Comments: Andrew, there are still a few technical issues here (especially about use of commas to set off phrases). But overall, this is better than the first version.

The major issue about the content I have is that you have some pretty strong claims here about the fact that Henrietta was taken advantage of by the doctors who collected the tissues from which the HeLa cell line was grown. You say that they were motivated by possible profits from selling the cells. Yet, I don’t see any of that motive in George Gey (at least in the way he is described in the book). He seems to me to be just a single-minded science “nerd” who is completely focused on his research, and in fact he ends up giving away all of the original samples to other scientists without generating anything in the way of monetary reward for himself.